OpenStudy (anonymous):

Suppose that you have a calculator that can only compute square roots. You need to calculate 7 to the 1/8 power. Remember that you do not have an exponent key. Explain how you could do this with this calculator. Explain why your method would work.

OpenStudy (anonymous):

\[a ^{1/x} = \sqrt[x]{a}\]\[7^{1/8}=\sqrt[8]{7}\]

OpenStudy (anonymous):

Thats what I thought but can you put that into a calculator?

OpenStudy (anonymous):

Yes. All you need is a \[\sqrt[n]{x}\] key.

OpenStudy (anonymous):

oh.. I dont think I have that button. Thank you again :)

OpenStudy (anonymous):

I'm sorry - It depends on what type of calculator you have. Most graphing and scientific calculators have that key...

OpenStudy (anonymous):

well I have the TI-83, Im probly just blind.. it took me 2 days to realize I had the LN button.. lol

OpenStudy (anonymous):

For the TI - 83: Click the MATH button (below ALPHA). Then select the key I showed above. I have the TI - 84 and the key is #5 in MATH.

OpenStudy (anonymous):

Ohh.. Isnt that snazzy, your the best :)

OpenStudy (anonymous):

My pleasure :)